A strong wind is causing a power line to oscillate.If the power line makes one complete up and down vibration every 5 seconds what is the power line's frequency?
1
Expert's answer
2018-05-27T08:38:08-0400
We can find the frequency of the power line from the formula:
f=1/T,
here, T is the time period of one complete up and down vibration.
Then, we get:
f=1/T=1/(5 s)=0.2 Hz.